\begin{thebibliography}{20}

%\cite{Bagger:2006sk}
\bibitem{Bagger:2006sk}
  J.~Bagger and N.~Lambert,
\emph{``Modeling multiple M2's,''}
  Phys.\ Rev.\  D {\bf 75} (2007) 045020
  [arXiv:hep-th/0611108].
  %%CITATION = PHRVA,D75,045020;%%


%\cite{Bagger:2007jr}
\bibitem{Bagger:2007jr}
  J.~Bagger and N.~Lambert,
\emph{``Gauge Symmetry and
Supersymmetry of Multiple M2-Branes,''}
  Phys.\ Rev.\  D {\bf 77} (2008) 065008
  [arXiv:0711.0955 [hep-th]].
  %%CITATION = PHRVA,D77,065008;%%


  %\cite{Bagger:2007vi}
\bibitem{Bagger:2007vi}
  J.~Bagger and N.~Lambert,
\emph{``Comments On Multiple
M2-branes,''}
  JHEP {\bf 0802} (2008) 105
  [arXiv:0712.3738 [hep-th]].
  %%CITATION = JHEPA,0802,105;%%


%\cite{Gustavsson:2008dy}
\bibitem{Gustavsson:2008dy}
  A.~Gustavsson,
\emph{``Selfdual strings and loop space
Nahm equations,''}
  JHEP {\bf 0804} (2008) 083
  [arXiv:0802.3456 [hep-th]].
  %%CITATION = JHEPA,0804,083

%\cite{Basu:2004ed}
\bibitem{Basu:2004ed}
  A.~Basu and J.~A.~Harvey,
\emph{``The M2-M5 brane system and a
generalized Nahm's equation,''}
  Nucl.\ Phys.\  B {\bf 713} (2005) 136
  [arXiv:hep-th/0412310].
  %%CITATION = NUPHA,B713,136;%%

%\cite{Schwarz:2004yj}
\bibitem{Schwarz:2004yj}
  J.~H.~Schwarz,
\emph{``Superconformal Chern-Simons
theories,''}
  JHEP {\bf 0411} (2004) 078
  [arXiv:hep-th/0411077].
  %%CITATION = JHEPA,0411,078;%%




%\cite{Mukhi:2008ux}
\bibitem{Mukhi:2008ux}
  S.~Mukhi and C.~Papageorgakis,
\emph{``M2 to D2,''}
  JHEP {\bf 0805} (2008) 085
  [arXiv:0803.3218 [hep-th]].
  %%CITATION = JHEPA,0805,085;%%

%\cite{Bandres:2008vf}
\bibitem{Bandres:2008vf}
  M.~A.~Bandres, A.~E.~Lipstein and J.~H.~Schwarz,
\emph{``N = 8 Superconformal
Chern--Simons Theories,''}
  JHEP {\bf 0805} (2008) 025
  [arXiv:0803.3242 [hep-th]].
  %%CITATION = JHEPA,0805,025;%%

%\cite{Berman:2008be}
\bibitem{Berman:2008be}
  D.~S.~Berman, L.~C.~Tadrowski and D.~C.~Thompson,
\emph{``Aspects of Multiple
Membranes,''}
  Nucl.\ Phys.\  B {\bf 802} (2008) 106
  [arXiv:0803.3611 [hep-th]].
  %%CITATION = NUPHA,B802,106;%%

%\cite{VanRaamsdonk:2008ft}
\bibitem{VanRaamsdonk:2008ft}
  M.~Van Raamsdonk,
\emph{``Comments on the Bagger-Lambert
theory and multiple M2-branes,''}
  JHEP {\bf 0805} (2008) 105
  [arXiv:0803.3803 [hep-th]].
  %%CITATION = JHEPA,0805,105;%%




%\cite{Morozov:2008cb}
\bibitem{Morozov:2008cb}
  A.~Morozov,
\emph{``On the Problem of Multiple M2
Branes,''}
  JHEP {\bf 0805} (2008) 076
  [arXiv:0804.0913 [hep-th]].
  %%CITATION = JHEPA,0805,076;%%

%\cite{Lambert:2008et}
\bibitem{Lambert:2008et}
  N.~Lambert and D.~Tong,
\emph{``Membranes on an Orbifold,''}
  arXiv:0804.1114 [hep-th].
  %%CITATION = ARXIV:0804.1114;%%

%\cite{Distler:2008mk}
\bibitem{Distler:2008mk}
  J.~Distler, S.~Mukhi, C.~Papageorgakis and M.~Van Raamsdonk,
\emph{``M2-branes on M-folds,''}
  JHEP {\bf 0805} (2008) 038
  [arXiv:0804.1256 [hep-th]].
  %%CITATION = JHEPA,0805,038;%%

%\cite{Gran:2008vi}
\bibitem{Gran:2008vi}
  U.~Gran, B.~E.~W.~Nilsson and C.~Petersson,
\emph{``On relating multiple M2 and
D2-branes,''}
  arXiv:0804.1784 [hep-th].
  %%CITATION = ARXIV:0804.1784;%%

%\cite{Ho:2008bn}
\bibitem{Ho:2008bn}
  P.~M.~Ho, R.~C.~Hou and Y.~Matsuo,
\emph{``Lie 3-Algebra and Multiple
M2-branes,''}
  JHEP {\bf 0806} (2008) 020
  [arXiv:0804.2110 [hep-th]].
  %%CITATION = JHEPA,0806,020;%%

%\cite{Gomis:2008cv}
\bibitem{Gomis:2008cv}
  J.~Gomis, A.~J.~Salim and F.~Passerini,
\emph{``Matrix Theory of Type IIB Plane
Wave from Membranes,''}
  arXiv:0804.2186 [hep-th].
  %%CITATION = ARXIV:0804.2186;%%

%\cite{Bergshoeff:2008cz}
\bibitem{Bergshoeff:2008cz}
  E.~A.~Bergshoeff, M.~de Roo and O.~Hohm,
\emph{``Multiple M2-branes and the
Embedding Tensor,''}
  Class.\ Quant.\ Grav.\  {\bf 25} (2008) 142001
  [arXiv:0804.2201 [hep-th]].
  %%CITATION = CQGRD,25,142001;%%

%\cite{Hosomichi:2008qk}
\bibitem{Hosomichi:2008qk}
  K.~Hosomichi, K.~M.~Lee and S.~Lee,
\emph{``Mass-Deformed Bagger-Lambert
Theory and its BPS Objects,''}
  arXiv:0804.2519 [hep-th].
  %%CITATION = ARXIV:0804.2519;%%

%\cite{Papadopoulos:2008sk}
\bibitem{Papadopoulos:2008sk}
  G.~Papadopoulos,
\emph{``M2-branes, 3-Lie Algebras and
Plucker relations,''}
  JHEP {\bf 0805} (2008) 054
  [arXiv:0804.2662 [hep-th]].
  %%CITATION = JHEPA,0805,054;%%

%\cite{Gauntlett:2008uf}
\bibitem{Gauntlett:2008uf}
  J.~P.~Gauntlett and J.~B.~Gutowski,
\emph{``Constraining Maximally
Supersymmetric Membrane Actions,''}
  arXiv:0804.3078 [hep-th].
  %%CITATION = ARXIV:0804.3078;%%


%\cite{Papadopoulos:2008gh}
\bibitem{Papadopoulos:2008gh}
  G.~Papadopoulos,
\emph{``On the structure of k-Lie
algebras,''}
  Class.\ Quant.\ Grav.\  {\bf 25} (2008) 142002
  [arXiv:0804.3567 [hep-th]].
  %%CITATION = CQGRD,25,142002;%%

%\cite{Ho:2008nn}
\bibitem{Ho:2008nn}
  P.~M.~Ho and Y.~Matsuo,
\emph{``M5 from M2,''}
  JHEP {\bf 0806} (2008) 105
  [arXiv:0804.3629 [hep-th]].
  %%CITATION = JHEPA,0806,105;%%

%\cite{Gomis:2008uv}
\bibitem{Gomis:2008uv}
  J.~Gomis, G.~Milanesi and J.~G.~Russo,
\emph{``Bagger-Lambert Theory for
General Lie Algebras,''}
  JHEP {\bf 0806} (2008) 075
  [arXiv:0805.1012 [hep-th]].
  %%CITATION = JHEPA,0806,075;%%

%\cite{Benvenuti:2008bt}
\bibitem{Benvenuti:2008bt}
  S.~Benvenuti, D.~Rodriguez-Gomez, E.~Tonni and H.~Verlinde,
\emph{``N=8 superconformal gauge
theories and M2 branes,''}
  arXiv:0805.1087 [hep-th].
  %%CITATION = ARXIV:0805.1087;%%

%\cite{Ho:2008ei}
\bibitem{Ho:2008ei}
  P.~M.~Ho, Y.~Imamura and Y.~Matsuo,
\emph{``M2 to D2 revisited,''}
  JHEP {\bf 0807} (2008) 003
  [arXiv:0805.1202 [hep-th]].
  %%CITATION = JHEPA,0807,003;%%


%\cite{Morozov:2008rc}
\bibitem{Morozov:2008rc}
  A.~Morozov,
\emph{``From Simplified BLG Action to
the First-Quantized M-Theory,''}
  arXiv:0805.1703 [hep-th].
  %%CITATION = ARXIV:0805.1703;%%

%\cite{Honma:2008un}
\bibitem{Honma:2008un}
  Y.~Honma, S.~Iso, Y.~Sumitomo and S.~Zhang,
\emph{``Janus field theories from
multiple M2 branes,''}
  arXiv:0805.1895 [hep-th].
  %%CITATION = ARXIV:0805.1895;%%

%\cite{Fuji:2008yj}
\bibitem{Fuji:2008yj}
  H.~Fuji, S.~Terashima and M.~Yamazaki,
\emph{``A New N=4 Membrane Action via
Orbifold,''}
  arXiv:0805.1997 [hep-th].
  %%CITATION = ARXIV:0805.1997;%%

%\cite{Ho:2008ve}
\bibitem{Ho:2008ve}
  P.~M.~Ho, Y.~Imamura, Y.~Matsuo and S.~Shiba,
\emph{``M5-brane in three-form flux and
multiple M2-branes,''}
  arXiv:0805.2898 [hep-th].
  %%CITATION = ARXIV:0805.2898;%%



%\cite{Song:2008bi}
\bibitem{Song:2008bi}
  Y.~Song,
\emph{``Mass Deformation of the
Multiple M2 Branes Theory,''}
  arXiv:0805.3193 [hep-th].
  %%CITATION = ARXIV:0805.3193;%%

%\cite{Jeon:2008bx}
\bibitem{Jeon:2008bx}
  I.~Jeon, J.~Kim, N.~Kim, S.~W.~Kim and J.~H.~Park,
\emph{``Classification of the BPS
states in Bagger-Lambert Theory,''}
  arXiv:0805.3236 [hep-th].
  %%CITATION = ARXIV:0805.3236;%%

%\cite{Li:2008ez}
\bibitem{Li:2008ez}
  M.~Li and T.~Wang,
\emph{``M2-branes Coupled to
Antisymmetric Fluxes,''}
  arXiv:0805.3427 [hep-th].
  %%CITATION = ARXIV:0805.3427;%%

%\cite{Hosomichi:2008jd}
\bibitem{Hosomichi:2008jd}
  K.~Hosomichi, K.~M.~Lee, S.~Lee, S.~Lee and J.~Park,
\emph{``N=4 Superconformal Chern-Simons
Theories with Hyper and Twisted Hyper
  Multiplets,''}
  arXiv:0805.3662 [hep-th].
  %%CITATION = ARXIV:0805.3662;%%



%\cite{Banerjee:2008pda}
\bibitem{Banerjee:2008pda}
  S.~Banerjee and A.~Sen,
\emph{``Interpreting the M2-brane
Action,''}
  arXiv:0805.3930 [hep-th].
  %%CITATION = ARXIV:0805.3930;%%

%\cite{Lin:2008qp}
\bibitem{Lin:2008qp}
  H.~Lin,
\emph{``Kac-Moody Extensions of
3-Algebras and M2-branes,''}
  arXiv:0805.4003 [hep-th].
  %%CITATION = ARXIV:0805.4003;%%

%\cite{DeMedeiros:2008zm}
\bibitem{DeMedeiros:2008zm}
  P.~De Medeiros, J.~Figueroa-O'Farril and E.~Mendez-Escobar,
\emph{``Lorentzian Lie 3-algebras and
their Bagger-Lambert moduli space,''}
  arXiv:0805.4363 [hep-th].
  %%CITATION = ARXIV:0805.4363;%%


%\cite{Gustavsson:2008bf}
\bibitem{Gustavsson:2008bf}
  A.~Gustavsson,
\emph{``One-loop corrections to
Bagger-Lambert theory,''}
  arXiv:0805.4443 [hep-th].
  %%CITATION = ARXIV:0805.4443;%%

%\cite{FigueroaO'Farrill:2008fz}
\bibitem{FigueroaO'Farrill:2008fz}
  J.~Figueroa-O'Farrill,
\emph{``Lorentzian Lie n-algebras,''}
  arXiv:0805.4760 [math.RT].
  %%CITATION = ARXIV:0805.4760;%%

%\cite{Bandres:2008kj}
\bibitem{Bandres:2008kj}
  M.~A.~Bandres, A.~E.~Lipstein and J.~H.~Schwarz,
\emph{``Ghost-Free Superconformal
Action for Multiple M2-Branes,''}
  arXiv:0806.0054 [hep-th].
  %%CITATION = ARXIV:0806.0054;%%

%\cite{Park:2008qe}
\bibitem{Park:2008qe}
  J.~H.~Park and C.~Sochichiu,
\emph{``Single M5 to multiple M2:
taking off the square root of
Nambu-Goto action,''}
  arXiv:0806.0335 [hep-th].
  %%CITATION = ARXIV:0806.0335;%%

%\cite{Passerini:2008qt}
\bibitem{Passerini:2008qt}
  F.~Passerini,
\emph{``M2-Brane Superalgebra from
Bagger-Lambert Theory,''}
  arXiv:0806.0363 [hep-th].
  %%CITATION = ARXIV:0806.0363;%%


%\cite{Gomis:2008be}
\bibitem{Gomis:2008be}
  J.~Gomis, D.~Rodriguez-Gomez, M.~Van Raamsdonk and H.~Verlinde,
\emph{``Supersymmetric Yang-Mills
Theory From Lorentzian
Three-Algebras,''}
  arXiv:0806.0738 [hep-th].
  %%CITATION = ARXIV:0806.0738;%%



%\cite{Aharony:2008ug}
\bibitem{Aharony:2008ug}
  O.~Aharony, O.~Bergman, D.~L.~Jafferis and J.~Maldacena,
\emph{``N=6 superconformal
Chern-Simons-matter theories, M2-branes
and their gravity duals,''}
  arXiv:0806.1218 [hep-th].
  %%CITATION = ARXIV:0806.1218;%%

%\cite{Ahn:2008ya}
\bibitem{Ahn:2008ya}
  C.~Ahn,
\emph{``Holographic Supergravity Dual
to Three Dimensional N=2 Gauge
Theory,''}
  arXiv:0806.1420 [hep-th].
  %%CITATION = ARXIV:0806.1420;%%

%\cite{Benna:2008zy}
\bibitem{Benna:2008zy}
  M.~Benna, I.~Klebanov, T.~Klose and M.~Smedback,
\emph{``Superconformal Chern-Simons
Theories and $AdS_4/CFT_3$
Correspondence,''}
  arXiv:0806.1519 [hep-th].
  %%CITATION = ARXIV:0806.1519;%%

%\cite{Ezhuthachan:2008ch}
\bibitem{Ezhuthachan:2008ch}
  B.~Ezhuthachan, S.~Mukhi and C.~Papageorgakis,
\emph{``D2 to D2,''}
  JHEP {\bf 0807} (2008) 041
  [arXiv:0806.1639 [hep-th]].
  %%CITATION = JHEPA,0807,041;%%

%\cite{Cecotti:2008qs}
\bibitem{Cecotti:2008qs}
  S.~Cecotti and A.~Sen,
\emph{``Coulomb Branch of the
Lorentzian Three Algebra Theory,''}
  arXiv:0806.1990 [hep-th].
  %%CITATION = ARXIV:0806.1990;%%



%\cite{Bergshoeff:2008ix}
\bibitem{Bergshoeff:2008ix}
  E.~A.~Bergshoeff, M.~de Roo, O.~Hohm and D.~Roest,
\emph{``Multiple Membranes from Gauged
Supergravity,''}
  arXiv:0806.2584 [hep-th].
  %%CITATION = ARXIV:0806.2584;%%

%\cite{deMedeiros:2008bf}
\bibitem{deMedeiros:2008bf}
  P.~de Medeiros, J.~Figueroa-O'Farrill and E.~Mendez-Escobar,
\emph{``Metric Lie 3-algebras in
Bagger-Lambert theory,''}
  arXiv:0806.3242 [hep-th].
  %%CITATION = ARXIV:0806.3242;%%

%\cite{Blau:2008bm}
\bibitem{Blau:2008bm}
  M.~Blau and M.~O'Loughlin,
\emph{``Multiple M2-Branes and Plane
Waves,''}
  arXiv:0806.3253 [hep-th].
  %%CITATION = ARXIV:0806.3253;%%

%\cite{Blau:2008bp}
\bibitem{Blau:2008bp}
  M.~Blau and M.~O'Loughlin,
\emph{``DLCQ and Plane Wave Matrix Big
Bang Models,''}
  arXiv:0806.3255 [hep-th].
  %%CITATION = ARXIV:0806.3255;%%

%\cite{Nishioka:2008gz}
\bibitem{Nishioka:2008gz}
  T.~Nishioka and T.~Takayanagi,
\emph{``On Type IIA Penrose Limit and
N=6 Chern-Simons Theories,''}
  arXiv:0806.3391 [hep-th].
  %%CITATION = ARXIV:0806.3391;%%


%\cite{Honma:2008jd}
\bibitem{Honma:2008jd}
  Y.~Honma, S.~Iso, Y.~Sumitomo and S.~Zhang,
\emph{``Scaling limit of N=6
superconformal Chern-Simons theories
and Lorentzian Bagger-Lambert
theories,''}
  arXiv:0806.3498 [hep-th].
  %%CITATION = ARXIV:0806.3498;%%



%\cite{Imamura:2008nn}
\bibitem{Imamura:2008nn}
  Y.~Imamura and K.~Kimura,
\emph{``Coulomb branch of generalized
ABJM models,''}
  arXiv:0806.3727 [hep-th].
  %%CITATION = ARXIV:0806.3727;%%



%\cite{Minahan:2008hf}
\bibitem{Minahan:2008hf}
  J.~A.~Minahan and K.~Zarembo,
\emph{``The Bethe ansatz for
superconformal Chern-Simons,''}
  arXiv:0806.3951 [hep-th].
  %%CITATION = ARXIV:0806.3951;%%


%\cite{Furuuchi:2008ki}
\bibitem{Furuuchi:2008ki}
  K.~Furuuchi, S.~Y.~Shih and T.~Takimi,
\emph{``M-Theory Superalgebra From
Multiple Membranes,''}
  arXiv:0806.4044 [hep-th].
  %%CITATION = ARXIV:0806.4044;%%


%\cite{Agarwal:2008rr}
\bibitem{Agarwal:2008rr}
  A.~Agarwal,
\emph{``Mass Deformations of Super
Yang-Mills Theories in D= 2+1, and
  Super-Membranes: A Note,''}
  arXiv:0806.4292 [hep-th].
  %%CITATION = ARXIV:0806.4292;%%

%\cite{Gaiotto:2008cg}
\bibitem{Gaiotto:2008cg}
  D.~Gaiotto, S.~Giombi and X.~Yin,
\emph{``Spin Chains in N=6
Superconformal Chern-Simons-Matter
Theory,''}
  arXiv:0806.4589 [hep-th].
  %%CITATION = ARXIV:0806.4589;%%

%\cite{Bandos:2008fr}
\bibitem{Bandos:2008fr}
  I.~A.~Bandos and P.~K.~Townsend,
\emph{``Light-cone M5 and multiple
M2-branes,''}
  arXiv:0806.4777 [hep-th].
  %%CITATION = ARXIV:0806.4777;%%

%\cite{Ahn:2008gda}
\bibitem{Ahn:2008gda}
  C.~Ahn,
\emph{``Towards Holographic Gravity
Dual of N=1Superconformal Chern-Simons
Gauge Theory,''}
  arXiv:0806.4807 [hep-th].
  %%CITATION = ARXIV:0806.4807;%%

%\cite{Bedford:2008hn}
\bibitem{Bedford:2008hn}
  J.~Bedford and D.~Berman,
\emph{``A note on Quantum Aspects of
Multiple Membranes,''}
  arXiv:0806.4900 [hep-th].
  %%CITATION = ARXIV:0806.4900;%%

%\cite{Arutyunov:2008if}
\bibitem{Arutyunov:2008if}
  G.~Arutyunov and S.~Frolov,
\emph{``Superstrings on $AdS_4 x CP^3$
as a Coset Sigma-model,''}
  arXiv:0806.4940 [hep-th].
  %%CITATION = ARXIV:0806.4940;%%

%\cite{Stefanski:2008ik}
\bibitem{Stefanski:2008ik}
  B.~.~j.~Stefanski,
\emph{``Green-Schwarz action for Type
IIA strings on $AdS_4\times CP^3$,''}
  arXiv:0806.4948 [hep-th].
  %%CITATION = ARXIV:0806.4948;%%

%\cite{Grignani:2008is}
\bibitem{Grignani:2008is}
  G.~Grignani, T.~Harmark and M.~Orselli,
\emph{``The $SU(2) x SU(2)$ sector in
the string dual of N=6 superconformal
Chern-Simons theory,''}
  arXiv:0806.4959 [hep-th].
  %%CITATION = ARXIV:0806.4959;%%

%\cite{Hosomichi:2008jb}
\bibitem{Hosomichi:2008jb}
  K.~Hosomichi, K.~M.~Lee, S.~Lee, S.~Lee and J.~Park,
 \emph{``N=5,6 Superconformal
 Chern-Simons Theories and M2-branes on
 Orbifolds,''}
  arXiv:0806.4977 [hep-th].
  %%CITATION = ARXIV:0806.4977;%%

%\cite{Fre:2008qc}
\bibitem{Fre:2008qc}
  P.~Fre and P.~A.~Grassi,
\emph{``Pure Spinor Formalism for
{Osp}(N|4) backgrounds,''}
  arXiv:0807.0044 [hep-th].
  %%CITATION = ARXIV:0807.0044;%%

%\cite{Bagger:2008se}
\bibitem{Bagger:2008se}
  J.~Bagger and N.~Lambert,
\emph{``Three-Algebras and N=6
Chern-Simons Gauge Theories,''}
  arXiv:0807.0163 [hep-th].
  %%CITATION = ARXIV:0807.0163;%%

%\cite{Terashima:2008sy}
\bibitem{Terashima:2008sy}
  S.~Terashima,
\emph{``On M5-branes in N=6 Membrane
Action,''}
  arXiv:0807.0197 [hep-th].
  %%CITATION = ARXIV:0807.0197;%%

%\cite{Grignani:2008te}
\bibitem{Grignani:2008te}
  G.~Grignani, T.~Harmark, M.~Orselli and G.~W.~Semenoff,
\emph{``Finite size Giant Magnons in
the string dual of N=6 superconformal
  Chern-Simons theory,''}
  arXiv:0807.0205 [hep-th].
  %%CITATION = ARXIV:0807.0205;%%

%\cite{Terashima:2008ba}
\bibitem{Terashima:2008ba}
  S.~Terashima and F.~Yagi,
\emph{``Orbifolding the Membrane
Action,''}
  arXiv:0807.0368 [hep-th].
  %%CITATION = ARXIV:0807.0368;%%

%\cite{Gromov:2008bz}
\bibitem{Gromov:2008bz}
  N.~Gromov and P.~Vieira,
\emph{``The $AdS4/CFT3$ algebraic
curve,''}
  arXiv:0807.0437 [hep-th].
  %%CITATION = ARXIV:0807.0437;%%

%\cite{Ahn:2008gd}
\bibitem{Ahn:2008gd}
  C.~Ahn and P.~Bozhilov,
\emph{``Finite-size effects of
Membranes on $AdS_4\times S_7$,''}
  arXiv:0807.0566 [hep-th].
  %%CITATION = ARXIV:0807.0566;%%

%\cite{Gromov:2008qe}
\bibitem{Gromov:2008qe}
  N.~Gromov and P.~Vieira,
\emph{``The all loop AdS4/CFT3 Bethe
ansatz,''}
  arXiv:0807.0777 [hep-th].
  %%CITATION = ARXIV:0807.0777;%%

%\cite{Chen:2008qq}
\bibitem{Chen:2008qq}
  B.~Chen and J.~B.~Wu,
\emph{``Semi-classical strings in
$AdS_4*CP^3$,''}
  arXiv:0807.0802 [hep-th].
  %%CITATION = ARXIV:0807.0802;%%






%\cite{Cherkis:2008qr}
\bibitem{Cherkis:2008qr}
  S.~Cherkis and C.~Saemann,
\emph{``Multiple M2-branes and
Generalized 3-Lie algebras,''}
  arXiv:0807.0808 [hep-th].
  %%CITATION = ARXIV:0807.0808;%%

%\cite{Chu:2008qv}
\bibitem{Chu:2008qv}
  C.~S.~Chu, P.~M.~Ho, Y.~Matsuo and S.~Shiba,
\emph{``Truncated Nambu-Poisson Bracket
and Entropy Formula for Multiple
  Membranes,''}
  arXiv:0807.0812 [hep-th].
  %%CITATION = ARXIV:0807.0812;%%

%\cite{Bandres:2008ry}
\bibitem{Bandres:2008ry}
  M.~A.~Bandres, A.~E.~Lipstein and J.~H.~Schwarz,
\emph{``Studies of the ABJM Theory in a
Formulation with Manifest SU(4)
R-Symmetry,''}
  arXiv:0807.0880 [hep-th].
  %%CITATION = ARXIV:0807.0880;%%

%\cite{Zhou:2008sa}
\bibitem{Zhou:2008sa}
  Y.~Zhou,
\emph{``Interpreting N M2 branes,''}
  arXiv:0807.0890 [hep-th].
  %%CITATION = ARXIV:0807.0890;%%

%\cite{Gomis:2008vc}
\bibitem{Gomis:2008vc}
  J.~Gomis, D.~Rodriguez-Gomez, M.~Van Raamsdonk and H.~Verlinde,
\emph{``A Massive Study of M2-brane
Proposals,''}
  arXiv:0807.1074 [hep-th].
  %%CITATION = ARXIV:0807.1074;%%

%\cite{Schnabl:2008wj}
\bibitem{Schnabl:2008wj}
  M.~Schnabl and Y.~Tachikawa,
\emph{``Classification of N=6
superconformal theories of ABJM
type,''}
  arXiv:0807.1102 [hep-th].
  %%CITATION = ARXIV:0807.1102;%%

%\cite{Li:2008ya}
\bibitem{Li:2008ya}
  T.~Li, Y.~Liu and D.~Xie,
 \emph{``Multiple D2-Brane Action from
 M2-Branes,''}
  arXiv:0807.1183 [hep-th].
  %%CITATION = ARXIV:0807.1183;%%

%\cite{Kim:2008gn}
\bibitem{Kim:2008gn}
  N.~Kim,
\emph{``How to put the Bagger-Lambert
theory on an orbifold : A derivation of
the ABJM model,''}
  arXiv:0807.1349 [hep-th].
  %%CITATION = ARXIV:0807.1349;%%

%\cite{Pang:2008hw}
\bibitem{Pang:2008hw}
  Y.~Pang and T.~Wang,
\emph{``From N M2's to N D2's,''}
  arXiv:0807.1444 [hep-th].
  %%CITATION = ARXIV:0807.1444;%%


%\cite{Garousi:2008ik}
\bibitem{Garousi:2008ik}
  M.~R.~Garousi, A.~Ghodsi and M.~Khosravi,
\emph{``On thermodynamics of N=6
superconformal Chern-Simons theories at
strong coupling,''}
  arXiv:0807.1478 [hep-th].
  %%CITATION = ARXIV:0807.1478;%%

%\cite{Hashimoto:2008iv}
\bibitem{Hashimoto:2008iv}
  A.~Hashimoto and P.~Ouyang,
\emph{``Supergravity dual of
Chern-Simons Yang-Mills theory with
N=6,8 superconformal IR fixed point,''}
  arXiv:0807.1500 [hep-th].
  %%CITATION = ARXIV:0807.1500;%%


%\cite{Astolfi:2008ji}
\bibitem{Astolfi:2008ji}
  D.~Astolfi, V.~G.~M.~Puletti, G.~Grignani, T.~Harmark and M.~Orselli,
\emph{``Finite-size corrections in the
$SU(2) x SU(2)$ sector of type IIA
string
  theory on $AdS_4 x CP^3$,''}
  arXiv:0807.1527 [hep-th].
  %%CITATION = ARXIV:0807.1527;%%

%\cite{AliAkbari:2008rm}
\bibitem{AliAkbari:2008rm}
  M.~Ali-Akbari, M.~M.~Sheikh-Jabbari and J.~Simon,
\emph{``Relaxed Three-Algebras: Their
Matrix Representations and Implications
for Multi M2-brane Theory,''}
  arXiv:0807.1570 [hep-th].
  %%CITATION = ARXIV:0807.1570;%%

%\cite{Ahn:2008aa}
\bibitem{Ahn:2008aa}
  C.~Ahn and R.~I.~Nepomechie,
\emph{``N=6 super Chern-Simons theory
S-matrix and all-loop Bethe ansatz
  equations,''}
  arXiv:0807.1924 [hep-th].
  %%CITATION = ARXIV:0807.1924;%%

%\cite{Hanaki:2008cu}
\bibitem{Hanaki:2008cu}
  K.~Hanaki and H.~Lin,
\emph{``M2-M5 Systems in N=6
Chern-Simons Theory,''}
  arXiv:0807.2074 [hep-th].
  %%CITATION = ARXIV:0807.2074;%%

%\cite{Verlinde:2008di}
\bibitem{Verlinde:2008di}
  H.~Verlinde,
\emph{``D2 or M2? A Note on Membrane
Scattering,''}
  arXiv:0807.2121 [hep-th].
  %%CITATION = ARXIV:0807.2121;%%

%\cite{Imamura:2008dt}
\bibitem{Imamura:2008dt}
  Y.~Imamura and K.~Kimura,
\emph{``N=4 Chern-Simons theories with
auxiliary vector multiplets,''}
  arXiv:0807.2144 [hep-th].
  %%CITATION = ARXIV:0807.2144;%%


%\cite{Lozano:1997ee}
\bibitem{Lozano:1997ee}
  Y.~Lozano,
\emph{``Eleven dimensions from the
massive D-2-brane,''}
  Phys.\ Lett.\  B {\bf 414} (1997) 52
  [arXiv:hep-th/9707011].
  %%CITATION = PHLTA,B414,52;%%

%\cite{Ortin:1997jh}
\bibitem{Ortin:1997jh}
  T.~Ortin,
\emph{``A note on the D-2-brane of the
massive type IIA theory and gauged
sigma models,''}
  Phys.\ Lett.\  B {\bf 415} (1997) 39
  [arXiv:hep-th/9707113].
  %%CITATION = PHLTA,B415,39;%%

%\cite{Romans:1985tz}
\bibitem{Romans:1985tz}
  L.~J.~Romans,
\emph{``Massive N=2a Supergravity In
Ten-Dimensions,''}
  Phys.\ Lett.\  B {\bf 169} (1986) 374.
  %%CITATION = PHLTA,B169,374;%%

%\cite{Bergshoeff:1996ui}
\bibitem{Bergshoeff:1996ui}
  E.~Bergshoeff, M.~de Roo, M.~B.~Green, G.~Papadopoulos and P.~K.~Townsend,
\emph{``Duality of Type II 7-branes and
8-branes,''}
  Nucl.\ Phys.\  B {\bf 470} (1996) 113
  [arXiv:hep-th/9601150].
  %%CITATION = NUPHA,B470,113;%%



  %\cite{Eyras:1998hn}
\bibitem{Eyras:1998hn}
  E.~Eyras, B.~Janssen and Y.~Lozano,
\emph{``5-branes, KK-monopoles and
T-duality,''}
  Nucl.\ Phys.\  B {\bf 531} (1998) 275
  [arXiv:hep-th/9806169].
  %%CITATION = NUPHA,B531,275;%%

%\cite{Bergshoeff:1998bs}
\bibitem{Bergshoeff:1998bs}
  E.~Bergshoeff and J.~P.~van der Schaar,
\emph{``On M-9-branes,''}
  Class.\ Quant.\ Grav.\  {\bf 16} (1999) 23
  [arXiv:hep-th/9806069].
  %%CITATION = CQGRD,16,23;%%

%\cite{Bergshoeff:1997ak}
\bibitem{Bergshoeff:1997ak}
  E.~Bergshoeff, Y.~Lozano and T.~Ortin,
\emph{``Massive branes,''}
  Nucl.\ Phys.\  B {\bf 518} (1998) 363
  [arXiv:hep-th/9712115].
  %%CITATION = NUPHA,B518,363;%%
%\cite{Bergshoeff:1997cf}
\bibitem{Bergshoeff:1997cf}
  E.~Bergshoeff, P.~M.~Cowdall and P.~K.~Townsend,
\emph{``Massive IIA supergravity from
the topologically massive D-2-brane,''}
  Phys.\ Lett.\  B {\bf 410} (1997) 13
  [arXiv:hep-th/9706094].
  %%CITATION = PHLTA,B410,13;%%

%\cite{Honma:2008ef}
\bibitem{Honma:2008ef}
  Y.~Honma, S.~Iso, Y.~Sumitomo, H.~Umetsu and S.~Zhang,
\emph{``Generalized Conformal Symmetry
and Recovery of SO(8) in Multiple M2
and D2 Branes,''}
  arXiv:0807.3825 [hep-th].
  %%CITATION = ARXIV:0807.3825;%%

%\cite{Bergshoeff:2008bh}
\bibitem{Bergshoeff:2008bh}
  E.~A.~Bergshoeff, O.~Hohm, D.~Roest, H.~Samtleben and E.~Sezgin,
\emph{``The Superconformal Gaugings in
Three Dimensions,''}
  arXiv:0807.2841 [hep-th].
  %%CITATION = ARXIV:0807.2841;%%

%\cite{Lee:2008ui}
\bibitem{Lee:2008ui}
  B.~H.~Lee, K.~L.~Panigrahi and C.~Park,
\emph{``Spiky Strings on AdS$_4 \times
{\bf CP}^3$,''}
  arXiv:0807.2559 [hep-th].
  %%CITATION = ARXIV:0807.2559;%%

%cite{Krishnan:2008zm}
\bibitem{Krishnan:2008zm} C.~Krishnan
and C.~Maccaferri, \emph{``Membranes on
Calibrations,''} arXiv:0805.3125
[hep-th].
%%CITATION = ARXIV:0805.3125;%%

\end{thebibliography}

\documentclass[hyper]{JHEP} 

\usepackage{epsfig}

\newcommand\fverb{\setbox\pippobox=\hbox\bgroup\verb}

\newcommand\fverbdo{\egroup\medskip\noindent%

            \fbox{\unhbox\pippobox}\ }

\newcommand\fverbit{\egroup\item[\fbox{\unhbox\pippobox}]}

\newbox\pippobox

\title{ D2 to
M2 Procedure for D2-Brane DBI
 Effective Action}
\author{by J. Kluso\v{n}\\
     Department of Theoretical Physics and Astrophysics\\
                   Faculty of Science, Masaryk University\\
Kotl\'{a}\v{r}sk\'{a} 2, 611 37, Brno\\
Czech Republic\\
    E-mail: \email{klu@physics.muni.cz}}

\preprint{\hepth{0807.4054}}

 \abstract{We apply
 the procedure
 that was suggested  in
[arXiv:0806.1639] to the case of
abelian D2-brane  Dirac-Born-Infeld
effective action  and discuss its
limitation. Then  we suggest
 an alternative form of this
  procedure that is based on an
   existence of
interpolating action proposed in
[hep-th/9707113,hep-th/9707011].}
\keywords{D-branes, M-branes}

\newcommand{\tB}{\tilde{B}}
\newcommand{\tD}{\tilde{D}}

\newcommand{\calg}{\mathcal{G}}
\newcommand{\calgi}{\left(\calg^{-1}\right)}

\newcommand{\hC}{\hat{C}}
\newcommand{\hg}{\hat{g}}
\newcommand{\hX}{\hat{X}}

\newcommand{\mF}{\mathcal{F}}

\newcommand{\tX}{\tilde{X}}

\newcommand{\hk}{\hat{k}}

\def \bAi{\left(\mathbf{A}^{-1}\right)}

\def \bA{\mathbf{A}}

\newcommand{\mL}{\mathcal{L}}

\begin{document}
\section{Introduction}\label{first}
It was proposed by Bagger and Lambert
in collection of very nice papers
\cite{Bagger:2006sk,Bagger:2007jr,Bagger:2007vi}
and independently by Gustavsson in
\cite{Gustavsson:2008dy} \footnote{For
related works, see
\cite{Mukhi:2008ux,Bandres:2008vf,Berman:2008be,VanRaamsdonk:2008ft,
Morozov:2008cb,Lambert:2008et,Distler:2008mk,Gran:2008vi,Ho:2008bn,
Gomis:2008cv,Bergshoeff:2008cz,Hosomichi:2008qk,Papadopoulos:2008sk,
Gauntlett:2008uf,Papadopoulos:2008gh,Ho:2008nn,Gomis:2008uv,Benvenuti:2008bt,
Ho:2008ei,Morozov:2008rc,Honma:2008un,Fuji:2008yj,Ho:2008ve,
Song:2008bi,Jeon:2008bx,Li:2008ez,Hosomichi:2008jd,Banerjee:2008pda,Lin:2008qp,
DeMedeiros:2008zm,Gustavsson:2008bf,FigueroaO'Farrill:2008fz,Bandres:2008kj,
Park:2008qe,Passerini:2008qt,Gomis:2008be,Ezhuthachan:2008ch,Cecotti:2008qs,
Bergshoeff:2008ix,deMedeiros:2008bf,Blau:2008bm,Blau:2008bp,Furuuchi:2008ki,
Agarwal:2008rr,Bandos:2008fr,Bedford:2008hn,Bagger:2008se,Terashima:2008sy,
Terashima:2008ba,Cherkis:2008qr,Chu:2008qv,Bandres:2008ry,Gomis:2008vc,
Schnabl:2008wj,Zhou:2008sa,Li:2008ya,Kim:2008gn,Pang:2008hw,
AliAkbari:2008rm,Hanaki:2008cu,Verlinde:2008di,Imamura:2008dt,Honma:2008ef,
Bergshoeff:2008bh,Krishnan:2008zm}.
 For study of supergravity duals of
these theories, see
\cite{Aharony:2008ug,Ahn:2008ya,Benna:2008zy,Nishioka:2008gz,
Honma:2008jd,Imamura:2008nn,Minahan:2008hf,Gaiotto:2008cg,Ahn:2008gda,
Arutyunov:2008if,Stefanski:2008ik,Grignani:2008is,Hosomichi:2008jb,
Fre:2008qc,Grignani:2008te,Gromov:2008bz,Ahn:2008gd,Gromov:2008qe,Chen:2008qq,
Garousi:2008ik,Hashimoto:2008iv,Astolfi:2008ji,Ahn:2008aa,Lee:2008ui}.}
following earlier works
\cite{Basu:2004ed,Schwarz:2004yj} that
a certain class of $N=8$
super-conformal theories in three
dimensions are potential candidates for
the world-volume description of
multiple M2-branes in M-theory. These
constructions are based on introducing
of an algebraic structure known as Lie
3-algebra that is needed for closure of
supersymmetry algebra. The metric
versions of the above theories fall
into two classes that depend on whether
the invariant bilinear form in
$3$-algebra space is positive definite
or indefinite. The original theories
proposed by Bagger-Lambert are
Euclidean theories with positive
definite bilinear form while more
recent proposals
\cite{Benvenuti:2008bt,Ho:2008ei}
contain bilinear form that is
indefinite  and these Lie 3-algebras
are known as Lorentzian 3-algebras.

It was claimed that the Lorentzian
3-algebra theories capture the
low-energy world-volume dynamics of
multiple parallel M2-branes. This model
has the required classical symmetries,
but has several unresolved problems. In
particular, the classical theory has
ghosts, $X_{\pm}$. Moreover, the
ghost-free formulation seems directly
equivalent to the non-conformal
D2-brane theory as was argued in
\cite{Bandres:2008kj,Gomis:2008be,Ezhuthachan:2008ch}.
Explicitly, it was shown very clearly
in \cite{Ezhuthachan:2008ch} how it is
possible-starting from $N=8$ SYM-
systematically and uniquely recovery
the theory
\cite{Bandres:2008kj,Gomis:2008be}.

Since the analysis presented in
\cite{Ezhuthachan:2008ch}
 is very nice and interesting it certainly deserves
further study. In fact, since $2+1$ dim
$N=8$ SYM theory describes low-energy
dynamics of $N$ D2-branes one can ask
the question whether it is possible to
extend this analysis
\cite{Ezhuthachan:2008ch}
 \footnote{In
what follows we call this analysis as
EMP procedure.}
 when we take
non-linear corrections into account.
 As the first  step in this
direction we try to apply  EMP
procedure
to the  case of single
Dirac-Born-Infeld (DBI) action for
D2-brane.

We start our analysis with the
remarkable form of $2+1$ dimensional
action that was proposed long ago in
\cite{Lozano:1997ee,Ortin:1997jh}. This
action is an interpolating action
 that-after appropriate integration of some
world-volume fields-either describes
D2-brane DBI effective action in
massive Type IIA supergravity or the
directly dimensional reduced gauged
M2-brane action.  We
 show that in linearized level this
action is equivalent to the abelian
form of the action given in
\cite{Ezhuthachan:2008ch} and hence can
be considered as the starting point for
non-linear generalization of EMP
procedure. On the other hand we argue
that naive application of EMP procedure
in this action leads to  a puzzle.
Explicitly, we argue that there is a
unique ground state of this new action
with infinite coupling constant. This
is different from we would expect since
M2 to D2-reduction is based on a
presumption that the vacuum expectation
value of $\left<X_+\right>$ can take
arbitrary constant value. Equivalently,
we would expect an infinite  number of
ground states that differ by vacuum
expectation values of $X_+$.

In order to resolve this problem we
suggest that the natural object for the
definition of non-linear EMP procedure
is gauged M2-brane action.
 More precisely, it is
well known that  in the case of IIA
supergravity it is possible to
introduce non-zero cosmological
constant proportional to $m^2$ with $m$
a mass parameter \cite{Romans:1985tz}.
Such backgrounds are essential for the
existence of D8-branes whose charge is
proportional to $m$
\cite{Bergshoeff:1996ui}. The action
for massive 11-dimensional theory has
the same contain as the massless one
\footnote{We use  the following notation for the hats.
Hats on target space fields
indicate that they are 11-dimensional.}
\begin{equation}
\hg_{MN} \ , \quad \hC_{MNK} \ , \quad
M,N,K=0,\dots, 11 \ .
\end{equation}
The action for these fields
 is manifestly  11-dimensional
Lorentz covariant but it does not
correspond to a proper 11-dimensional
theory because, in order to write down
the action, we need to introduce an
auxiliary non-dynamic vector field
$\hk^M$ such that the Lie derivatives
of the metric and 3-form potential with
respect to it are zero:
\begin{equation}
\mL_{\hk}\hg_{MN}=0 \ , \quad
\mL_{\hk}\hC_{MNL}=0 \ .
\end{equation}
An existence of this Killing vector is
crucial for definition of massive
M2-brane. In fact, the
 world-volume theory of
massive branes \footnote{These
branes-that propagate in the background
with non-zero cosmological constant-are
called as "massive branes" as opposed
to branes that propagate in a
background with zero mass parameter. It
is clear that all these  branes are
massive in the sense that their
physical mass is nonzero.} was
extensively studied in the past, for
example
\cite{Eyras:1998hn,Bergshoeff:1998bs,Bergshoeff:1997ak,Bergshoeff:1997cf}.
These actions  have as a common
characteristic that they are gauged
sigma models. The gauged isometry is
the same as an isometry that is needed
in order to define the massive
11-dimensional supergravity theory. For
example, the original un-gauged
M2-brane action is the same object as
in the  massless theory, i.e. the
corresponding massless  M2-brane.

Let us again return to generalized EMP
procedure. We argue that it can be
naturally applied for gauged M2-brane
action.  As opposite to the original
EMP procedure, where  the Yang-Mills
coupling constant  vector $g_{YM}^I$ is
replaced with a dynamical field $X_+^I$
we replace the constant Killing vector
 $\hk^M$ with dynamical field
$\hX_+^M$. Then we can easily find
manifestly covariant form of the
generalized action with infinite number
of ground states that differ by vacuum
expectation values of $\hX^M_+$.

It is remarkable that the gauged
isometry that  appears in massive
M2-brane action is related to the gauge
symmetry introduced in
\cite{Ezhuthachan:2008ch}. We hope that
this observation will allow  to find
new geometrical interpretations of
gauge symmetries that were introduced
in
\cite{Bandres:2008kj,Gomis:2008be,Ezhuthachan:2008ch}.

The organization of this paper is as
follows. In the next section
(\ref{second}) we introduce the
interpolating D2-brane action and we
argue that after appropriate
redefinition of world-volume fields it
agrees with the abelian version of
D2-brane action introduced in
\cite{Ezhuthachan:2008ch}.  In section
(\ref{third}) we apply EMP prescription
for gauged M2-brane action and we find
covariant and non-linear version of
M2-brane action that has all desired
properties. In conclusion
(\ref{fourth}) we outline our results
and suggest possible extension of our
work. Finally, in Appendix
\ref{Appendix} we explicitly show that
the dimensional reduction in gauged
M2-brane action leads to the
interpolating action introduced in
section (\ref{second}).

\section{D2-Brane Action}\label{second}
We start with the action that was proposed
in \cite{Lozano:1997ee,Ortin:1997jh}
\begin{eqnarray}\label{Ortinact}
S[X^m,X,&V_\mu &,B_\mu]=
\nonumber \\
&=&-\tau_{M2}
\int d^3\xi  e^{-\Phi}\sqrt{-\det
[g_{\mu\nu}+e^{2\Phi}
F_\mu F_\nu]}+\nonumber \\
&+&\frac{\tau_{M2}}{3!} \int d^3\xi
\epsilon^{\mu\nu\rho}
[C^{(3)}_{\mu\nu\rho}+6\pi\alpha'D_\mu
X\mF_{\nu\rho} +6m
(\pi\alpha')^2V_\mu\partial_{\nu}V_{\rho}]
\ ,
\nonumber \\
\end{eqnarray}
where
\begin{eqnarray}
F_\mu &=& D_\mu X+C^{(1)}_\mu \ ,
\nonumber \\
D_\mu X&=&\partial_\mu X+B_\mu \ ,
\nonumber \\
\mF_{\mu\nu}&=&\partial_\mu
V_\nu-\partial_\nu V_\mu-
\frac{1}{2\pi\alpha'}b_{\mu\nu}
\nonumber \\
\end{eqnarray}
and where
\begin{eqnarray}
g_{\mu\nu}&=&\partial_\mu
X^m\partial_\nu X^n g_{mn} \ , \quad
b_{\mu\nu}=b_{mn}\partial_\mu
X^m\partial_\nu X^n \ , \nonumber \\
C_{\mu\nu\rho}^{(3)}&=&C_{mnk}
\partial_\mu X^m\partial_\nu X^n\partial_\rho X^k \ , \quad
C^{(1)}_\mu=C_m^{(1)}\partial_\mu X^m \
, \nonumber \\
\end{eqnarray}
where $g_{mn},b_{mn}$ are space-time
metric and NS two form field
respectively and where
$C^{(3)}_{mnk},C^{(1)}_m$ are
Ramond-Ramond three and one forms
respectively. Further,
 $X^m \ , m,n=0,\dots,9$ are
world-volume modes that describe embedding
of D2-brane in the  target space-time.
Finally, $\tau_{M2}$ is D2-brane
tension defined as
$\tau_{M2}=\frac{1}{l_s^3}$.

 The action (\ref{Ortinact}) contains
extra fields as opposite to usual DBI
action for D2-brane. Firstly, if we
integrate out $B_\mu$ we obtain
\begin{eqnarray}\label{eqB}
-\frac{e^{\Phi}\sqrt{-g}g^{\mu\nu}F_\nu}{
\sqrt{1+e^{2\Phi}g^{\mu\nu}F_\mu
F_\nu}}+
\pi\alpha'\epsilon^{\mu\nu\rho}
\mF_{\nu\rho}=0 \ ,
\end{eqnarray}
where  we used the
fact that
\begin{eqnarray}
\sqrt{-\det[g_{\mu\nu}+e^{2\Phi}F_\mu
F_\nu]}
=\sqrt{-\det g}\sqrt{1+e^{2\Phi}
g^{\mu\nu}F_\mu F_\nu}\ .
\end{eqnarray}
Then if we insert (\ref{eqB}) into
(\ref{Ortinact}) we obtain an action in the
form
\begin{eqnarray}\label{OrtinactDBI}
S= -\tau_{M2} \int d^3\xi e^{-\Phi}
\sqrt{-\det
[g_{\mu\nu}+2\pi\alpha'\mF_{\mu\nu}]}
+\nonumber \\
+\frac{\tau_{M2}}{3!} \int d^3\xi
\epsilon^{\mu\nu\rho}
(C^{(3)}_{\mu\nu\rho}-6\pi\alpha'C_{\mu}^{(1)}\mF_{\nu\rho}
+6m
(\pi\alpha')^2V_\mu\partial_{\nu}V_{\rho})
\
\nonumber \\
\end{eqnarray}
that is standard form of D2-brane in massive
Type IIA background and that reduces  to
the massless Type IIA background when
$m=0$.

In order to see that the action
(\ref{OrtinactDBI}) is related to
abelian reduction of the action given
in \cite{Ezhuthachan:2008ch}  we take
 following background:
\begin{equation}
g_{mn}=\eta_{mn}  \ , \quad
 \Phi=\Phi_0=\mathrm{const} \ , \quad
C^{(1)}_m=C^{(3)}_{mnk}=0 \ .
\end{equation}
Further, let us impose static gauge
\begin{equation}
X^\mu=\xi^\mu \ , \quad \mu=0,1,2
\end{equation}
 so that
\begin{equation}
g_{\mu\nu}=\eta_{\mu\nu}+\delta_{ij}\partial_\mu
X^i\partial_\nu X^j \ , \quad i,j=3,\dots,9
\ .
\end{equation}
Then in the quadratic approximation the
action (\ref{OrtinactDBI}) takes the form
\begin{eqnarray}\label{OrtinactDBIq}
S[X^m,X,V_\mu,B_\mu]&=&-\tau_{M2}\int d^3\xi e^{-\Phi_0}
\sqrt{-\eta}-\tau_{M2}\int d^3 \xi
\sqrt{-\eta}
[\frac{1}{2}e^{-\Phi_0}\eta^{\mu\nu}
\delta_{ij}\partial_\mu X^i
\partial_\nu X^j+\nonumber \\
&+&\frac{e^{\Phi_0}}{2}
\eta^{\mu\nu}F_\mu F_\nu]+
\int d^3\xi \epsilon^{\mu\nu\rho}
(\pi\alpha' \tau_{M2})D_\mu X \mF_{\nu\rho} \ .
\nonumber \\
\end{eqnarray}
As the next step we introduce
the gauge theory  coupling constant
through the standard relations
\begin{eqnarray}
e^{-\Phi_0}l_s^4\tau_{M2}=\frac{1}{g^2_{YM}}
\ , \quad  \left(\tau_{M2}=
\frac{1}{l_s^3} \ , \quad
2\pi\alpha'=l_s\right) \  \quad
\nonumber \\
\end{eqnarray}
so that after rescaling
\begin{equation}\label{rescaling}
\sqrt{\tau_{M2}} e^{\Phi_0/2}X=\tX \ ,
\quad \sqrt{\tau_{M2}} e^{-\Phi_0/2}X^i=
\tX^i \ , \quad
\frac{1}{l_s^{5/2}}B_\mu=\tB_\mu \ , \quad
l_s^{3/2}\mF_{\mu\nu}=\tilde{\mF}_{\mu\nu}
\end{equation}
the action (\ref{OrtinactDBIq})
 takes the form
\begin{eqnarray}
S[\tX^m,\tX,\tilde{V}_\mu,\tB]&=&-\frac{1}{g^2_{YM}l_s^2}\int d^3\xi
\sqrt{-\eta}- \int d^3 \xi \sqrt{-\eta}
[\frac{1}{2}\eta^{\mu\nu}
\delta_{ij}\partial_\mu \tX^i
\partial_\nu \tX^j+\nonumber \\
&+&\frac{1}{2} \eta^{\mu\nu}
(\partial_\mu \tX+g_{YM}\tB_\mu)
(\partial_\nu
\tX+g_{YM}\tB_\nu)]+\nonumber \\
&+& \int
d^3\xi \epsilon^{\mu\nu\rho}
[\frac{1}{2} \tB_\mu \mF_{\nu\rho} +
\frac{1}{2l_s^{3/2}g_{YM}}
\partial_\mu \tX \tilde{\mF}_{\nu\rho}
]
\nonumber \\
\end{eqnarray}
that has the same form as the abelian
form of the action given in
\cite{Ezhuthachan:2008ch}
\footnote{This is true up to total
derivative term since $\int d^3\xi
\epsilon^{\mu\nu\rho}
\partial_\mu \tX \tilde{\mF}_{\nu\rho}=
\int d^3\xi \partial_\mu [\epsilon^{\mu\nu\rho}
\tX \tilde{\mF}_{\nu\rho}]-
\int d^3\xi \tX\partial_\mu(\epsilon^{\mu\nu\rho}
\tilde{\mF}_{\nu\rho})=
\int d^3\xi \partial (\dots)$.}.
Motivated by this result we perform
the rescaling (\ref{rescaling}) in
the action (\ref{OrtinactDBI}) and
we obtain
\begin{eqnarray}\label{OrtinactDBI2}
S[\tX^m,\tX,\tilde{V}_\mu,\tB_\mu]&=&
-\int d^3\xi \sqrt{-\det\bA_{\mu\nu}} +
\int d^3\xi \epsilon^{\mu\nu\rho}
\frac{1}{2} \tB_\mu
\tilde{\mF_{\nu\rho}} \ ,\nonumber \\
\bA_{\mu\nu}&=&
\frac{1}{l_s^{8/3}g_{YM}^{4/3}}\eta_{\mu\nu}+g_{YM}^{2/3}
l_s^{4/3}
\partial_\mu \tX^i\partial_\nu
\tX^j\delta_{ij}+\nonumber \\
&+&l_s^{4/3}g_{YM}^{2/3}(\partial_\mu \tX+
g_{YM}\tB_\mu)(\partial_\nu\tX+g_{YM}\tB_\nu)
 \ ,
\nonumber \\
\end{eqnarray}
where we ignored term that contributes
to the action as total derivative. Now
we are ready to apply EMP procedure for
(\ref{OrtinactDBI2}). We introduce
$8$-dimensional vector $g_{YM}^I$ as
$g_{YM}^I=(\overbrace{0,\dots,0}^{7},g_{YM}),I=1,\dots,8,
$ and "covariant derivative
$\tD$"
\begin{equation}
\tD_\mu \tX^i=\partial_\mu
\tX^i+g_{YM}^i \tB_\mu \ , \quad  \tD_\mu
\tX=\partial_\mu \tX+g_{YM}\tB_\mu \ .
\end{equation}
 Further, we rewrite
 $g_{YM}^2$ in manifest $SO(8)$ covariant manner
 as $g^2_{YM}=g^I_{YM}g^J_{YM}\delta_{IJ}=
|g_{YM}|^2$
and then we replace
vector $g_{YM}^I$ with dynamical field
$X^I_+$ so that the action
(\ref{OrtinactDBI2})
 takes the
form
\begin{eqnarray}\label{OrtinactDBI4}
&S&[\tX^I,\tX_+^I,\tilde{V}_\mu,\tB_\mu,C^\mu_{I}]=
\nonumber \\
&=&-\int d^3\xi\left( \sqrt{-\det
\bA_{\mu\nu}} +\epsilon^{\mu\nu\rho}
\frac{1}{2} \tB_\mu
\tilde{\mF_{\nu\rho}}
+ C^\mu_I \partial_\mu \tX^I_+\right) \
,
\nonumber \\
\bA_{\mu\nu}&=&\frac{1}{l_s^{8/3}
(X^I_+X^J_+\delta_{IJ})^{2/3}}\eta_{\mu\nu}+(
X^I_+X_+^J\delta_{IJ})^{1/3}
l_s^{4/3}\tD_\mu \tX^I\tD_\nu
\tX^J\delta_{IJ} \ ,  \nonumber \\
\end{eqnarray}
where we introduced auxiliary
field $C^\mu_I$ that renders $\tX^I_+$
non-dynamical.

Let us now analyze some  properties of
the action (\ref{OrtinactDBI4}). We are
mainly interested in the study of the
ground state of this theory that has to
solve the equations of motion that
follow from the action
(\ref{OrtinactDBI4}). We presume that
the ground state is characterized by
following configuration of the
world-volume fields
\begin{equation}\label{grst}
\tX^8_+=v=\mathrm{const} \ , \quad
\tD_\mu \tX^I=0 \ , \quad \tB_\mu=0 \ ,
\quad  \tilde{\mF}_{\mu\nu}=0 \ .
\end{equation}
Firstly, the equation of motion for
$C^\mu_I$ takes the form
\begin{equation}
\partial_\mu \tX^I_+=0
\end{equation}
that is clearly obeyed by the ansatz
(\ref{grst}).
On the other hand the equation of
motion for $X^I_+$ takes the form
\begin{eqnarray}\label{eqX+}
-\frac{1}{2}\frac{\delta \bA_{\mu\nu}}
{\delta
X^I_+}\bAi^{\nu\mu}\sqrt{-\det\bA}
-\epsilon^{\mu\nu\rho}
\frac{1}{2l_s^{1/2}(X^I_+X^J_+\delta_{IJ})^{3/2}}
\partial_\mu \tX
\tilde{\mF}_{\nu\rho}=0 \ .
\nonumber \\
\end{eqnarray}
Since  for the ansatz (\ref{grst}) the
matrix $\bA_{\mu\nu}$ is equal to
\begin{equation}
\bA_{\mu\nu}=\frac{1}{l_s^{8/3}v^{4/3}}\eta_{\mu\nu}
\end{equation}
the equation (\ref{eqX+})
implies
\begin{equation}
\frac{1}{v^{3}}=0  \ .
\end{equation}
In other words the ground state
corresponds to the point $v\rightarrow
\infty$ that implies that there is
unique ground state of the theory. As
we argued in introduction this is not
the same what we want since we would
like to have a theory with infinite
number of ground states that differ by
vacuum expectation values of $\tX_+$.
In order to find solution of this
problem we suggest
 an alternative
procedure how to introduce $X_+^I$ as a
new dynamical variable. In the next
section we present such an alternative
procedure that is based on the fact
that the action (\ref{OrtinactDBI}) can
be considered as dimensional reduction
of massive M2-brane.
\section{Gauged Theory for
M2-brane}\label{third}
 Let us again
consider the action (\ref{OrtinactDBI})
and determine the equations of motion
for $V_\mu$
\begin{equation}\label{eqV}
 \pi m\alpha' \epsilon^{\mu\nu\rho}
(\partial_\nu V_\rho-\partial_\rho
V_\mu)+\epsilon^{\mu\nu\rho}
(\partial_\nu B_\rho-\partial_\rho
B_\nu)=0 \ .
\end{equation}
Inserting (\ref{eqV}) back to
(\ref{OrtinactDBI}) we obtain the
action in the form
\begin{eqnarray}\label{OrtinactDBIdr}
S=-\tau_{M2}\int d^3\xi
\sqrt{-\det (e^{-2/3\Phi}
g_{\mu\nu}+e^{4/3\Phi}F_\mu F_\nu)}+
 \nonumber \\
+\frac{\tau_{M2}}{6} \int d^3\xi
\epsilon^{\mu\nu\rho}
(C_{\mu\nu\rho}^{(3)}-3D_\mu X
B^{(1)}_{\nu\rho} +\frac{6}{m}B_\mu
\partial_\nu B_\rho) \ .
\nonumber \\
\end{eqnarray}
As was shown in \cite{Ortin:1997jh}
(and reviewed in appendix
\ref{Appendix}) this action is very
close to the action that one gets by
direct dimensional reduction of the
massive M2-brane that is also known as
gauged M2-brane action. This action can
be defined in the background with
Killing vector isometry $\hk^M(\hX)$.
Then the
  gauged M2-bane action
takes the form \cite{Ortin:1997jh}
\begin{eqnarray}\label{M2gauged}
S&=&-\tau_{M2}\int d^3\xi \sqrt{-\det
D_\mu \hX^M D_\nu \hX^N \hg_{MN}}
+\nonumber  \nonumber \\
&+& \tau_{M2}\int d^3\xi
\epsilon^{\mu\nu\rho} [D_\mu \hX^M
D_\nu \hX^N D_\rho \hX^K
\hC_{MNK}-\frac{6}{m}B_\mu \partial_\nu
B_\rho] \ ,
\nonumber \\
\end{eqnarray}
where the covariant derivative $D_\mu$
is defined as
\begin{equation}
D_\mu \hX^M=\partial_\mu \hX^M +B_\mu
\hk^M(\hX) \ ,
\end{equation}
where $B_\mu$ is world-volume gauge
field related to the Killing gauge
isometry. To clarify meaning of this
gauged form of the action let us
consider following transformation
\begin{equation}\label{deltahx}
\delta_\eta\hX^M(\xi)=
\hX'^M(\xi)-\hX^M(\xi)=
\eta(\xi)\hk^M(\hX) \ ,
\end{equation}
where $\eta(\xi)$ is a parameter of
gauge transformations. This
transformation immediately implies
following transformation rules of
background fields
\begin{eqnarray}\label{deltag}
\delta_\eta \hg_{MN}=\eta \hk^K
\partial_K \hg_{MN} \ , \quad
\delta_\eta \hC_{KMN}=\eta \hk^L\partial_L
\hC_{KMN} \ , \quad
\delta_\eta \hk^K=\eta \hk^L\partial_L \hk^K
\nonumber \\
\end{eqnarray}
and transformation of covariant
derivative
\begin{eqnarray}
\delta_\eta D_\mu \hX^M=
\eta D_\mu
\hX^L\partial_L\hk^M \ ,
\end{eqnarray}
where we postulate following
transformation rule for gauge field
$B_\mu$
\begin{equation}
\delta B_\mu=-\partial_\mu \eta \ .
\end{equation}
Then
\begin{eqnarray}
\delta (D_\mu \hX^M D_\nu \hX^N
\hg_{MN})= \eta  D_\mu \hX^M(
\partial_M \hk^L \hg_{LN}+
\hg_{ML}\partial_N \hk^L+
\partial_L \hg_{MN})D_\nu \hX^N=0
\nonumber \\
\end{eqnarray}
since
\begin{equation}
\mL_{\hk}\hg_{MN}=0 \ .
\end{equation}
In the same way we obtain that
\begin{equation}
\delta_\eta (D_\mu \hX^M D_\nu \hX^N
D_\rho \hX^K \hC_{MNK})=0 \ , \quad
\mL_{\hk}\hC_{MNK}=0
\end{equation}
hence we see that the action is
invariant under  transformations
(\ref{deltahx}) and (\ref{deltag}).

Having clarified the fact that the
D2-brane action (\ref{OrtinactDBI}) is
related to the gauged M2-brane action
we now introduce  modified EMP
procedure to the action
(\ref{M2gauged}). As the first step in
our construction we will presume an
existence of  adapted system of
coordinates where $\hk^M=const$. This
is always possible to achieve in flat
background
$\hg_{MN}=\eta_{MN},\hC_{MNK}=0$.
Further, in analogy with EMP
prescription, we replace constant
$\hk^M$ with  dynamical field $\hX^M_+$
and add to the action term
$\frac{1}{2}C^\mu_M
\partial_\mu \hX^M_+$ to render this field
non-dynamical.

Further  we rewrite the Wess-Zumino
term in (\ref{M2gauged}) as
\begin{equation}\label{Vin}
\pi\alpha'\epsilon^{\mu\nu\rho} B_\mu
\mF_{\nu\rho}+ (\pi\alpha')^2m
\epsilon^{\mu\nu\rho}V_\mu\partial_\nu
V_\rho \ .
\end{equation}
In fact it is easy to see that now
the equation of motion for $V_\mu$
that follow from (\ref{Vin}) implies
\begin{equation}
\frac{1}{m\pi\alpha'}
 (\partial_\mu
B_\nu-\partial_\nu B_\mu)= -
(\partial_\mu V_\nu-\partial_\nu V_\mu)
\end{equation}
and hence when we insert it back to (\ref{Vin})
 we obtain the last term in (\ref{M2gauged}). Note
also that this expression is invariant
under $\eta$ transformations (up to
total derivative) since
\begin{equation}
\delta (\epsilon^{\mu\nu\rho} B_\mu
\mF_{\nu\rho})= \epsilon^{\mu\nu\rho}
\partial_\mu \eta \mF_{\nu\rho}=
-\eta\epsilon^{\mu\nu\rho}
\partial_\mu\partial_\nu A_\rho+
\eta\epsilon^{\mu\nu\rho}
\partial_\mu\partial_\rho A_\nu=0 \ .
\end{equation}
In summary we derive the action in the
form
\begin{eqnarray}
S&=&-\tau_{M2}\int d^3\xi [\sqrt{-\det
\calg_{\mu\nu}}+\frac{1}{2}\sqrt{-\det
\calg_{\mu\nu}}C_\nu^N\eta_{NM}
\calgi^{\nu\mu}\partial_\mu\hX^M_+
+\nonumber  \\
&+& \frac{\tau_{M2}}{3!}\int d^3\xi
\epsilon^{\mu\nu\rho} [6\pi\alpha'
B_\mu \mF_{\nu\rho}+6m(\pi\alpha')^2
V_\mu
\partial_\nu V_\rho] \ ,
\nonumber \\
\end{eqnarray}
where we  added term $\frac{1}{2}
C^M_\mu \calgi^{\mu\nu}\eta_{MN}\partial_\nu
\hX^N_+$ that renders $\hX_+^M$
constant on-shell and
where we also introduced
"generalized metric" $\calg_{\mu\nu}$
\begin{equation}
\calg_{\mu\nu}= (\partial_\mu
\hX^M+B_\mu \hX^M_+)
\eta_{MN}(\partial_\nu \hX^N+ B_\nu
\hX^N_+) \ .
\end{equation}

Following \cite{Ezhuthachan:2008ch}
 we introduce   field $\hX^M_-$ and add to
the action an expression
\[\frac{1}{2}\sqrt{-\det\calg}
\partial_\mu \tX^M_-\eta_{MN}\calgi^{\mu\nu}
\partial_\nu \tX^N_+ \]
in order the action will be invariant
under  additional shift symmetry
\begin{equation}\label{deltash}
\delta C^M_\mu=\partial_\mu\lambda^M
 \ , \quad  \delta \hX^M_{-}=
\lambda^M \ .
\end{equation}
Then the final form of the action
 takes the form
\begin{eqnarray}
S&=&-\tau_{M2}\int d^3\xi \sqrt{-\det
\calg_{\mu\nu}}\left(1+\right.\nonumber \\
&+& \left.\frac{1}{2}(C_{\mu}^M
-\partial_\mu\hX_{-}^M)
\eta_{MN}
 \calgi^{\mu\nu}
\partial_\nu
\hX^M_+\right)
+\nonumber  \\
&+& \frac{\tau_{M2}}{3!}\int d^3\xi
\epsilon^{\mu\nu\rho} [6\pi\alpha'
B_\mu \mF_{\nu\rho}+6m(\pi\alpha')^2
V_\mu
\partial_\nu V_\rho]
\nonumber \\
\end{eqnarray}
Note also that in order to achieve that
 $\hX_+^M$ is constant on-shell and
that the action possesses additional
shift symmetry we can consider more
general form of the action
\begin{eqnarray}
S&=&-\tau_{M2}\int d^3\xi \sqrt{-\det
\calg_{\mu\nu}}
\sqrt{1+(C_\mu^M-\partial_\mu \hX_-^M)
\eta_{MN}\calgi^{\mu\nu}\partial_\nu \hX^N_+}
-\nonumber  \\
&-& \frac{\tau_{M2}}{3!}\int d^3\xi
\epsilon^{\mu\nu\rho} [6\pi\alpha'
B_\mu \mF_{\nu\rho}+6m(\pi\alpha')^2
V_\mu
\partial_\nu V_\rho]
\nonumber \\
\end{eqnarray}
that can be finally written in a
suggestive form as
\begin{eqnarray}\label{M2final}
S&=&-\tau_{M2}\int d^3\xi \sqrt{-\det
\bA_{\mu\nu}}
-\nonumber \\
&-& \frac{\tau_{M2}}{3!}\int d^3\xi
\epsilon^{\mu\nu\rho} [6\pi\alpha'
B_\mu \mF_{\nu\rho}+6m(\pi\alpha')^2
V_\mu
\partial_\nu V_\rho]
\nonumber \\
\bA_{\mu\nu}&=&
\calg_{\mu\nu}+(C_\mu^M-\partial_\mu \hX^M_-)
\eta_{MN}\partial_\nu \hX^N_+ \ .
\nonumber \\
\end{eqnarray}
Let us now study properties of the
action (\ref{M2final}).  Clearly it
 is invariant
under shift symmetry (\ref{deltash}).
Further, the variation of this action with respect to
$C^M_\mu$ implies
\begin{equation}
\eta_{NM}\partial_\nu \hX^M_+
\bAi^{\nu\mu}\sqrt{-\det\bA}=0
\end{equation}
that  implies $\partial_\nu \hX^M_+=0$.
Let us again presume the ground state
of the theory in the form
\begin{equation}\label{anstM}
B_\mu=\hX^M_-=C^M_\mu=V_\mu=0 \ , \quad
\hX^M_+=v^M \ .
\end{equation}
It is easy to see that the equations
of motion for $\hX^M,B_\mu,C^M_\mu$ and
$\hX^M_-$ are obeyed for this ansatz. Finally,
the problematic equation of motion for
$\hX^M_+$ takes the form
\begin{eqnarray}
& & B_\mu \eta_{MN}(\partial_\nu
\hX^N+B_\nu \hX^N_+)\bAi^{\nu\mu}
\sqrt{-\det\bA}+\nonumber \\
&+&\frac{1}{2}
\partial_\mu[\eta_{MN}\partial_\nu
\hX^N_-\bAi^{\nu\mu}\sqrt{-\det\bA}]=0
\nonumber \\
\end{eqnarray}
that is clearly solved by (\ref{anstM})
for any $v^M$. Finally, let us impose
the static gauge in the following form
\begin{eqnarray}
\hX^\mu=\xi^\mu \ , \quad  \mu \ , \nu=0,1,2 \nonumber \\
C_\mu^\nu=0 \ , \quad
\hX^\mu_+=\hX^\mu_-=0
\nonumber \\
\end{eqnarray}
so that
the matrix $\bA_{\mu\nu}$ takes the form
\begin{eqnarray}
\bA_{\mu\nu}&=&
\eta_{\mu\nu}+(\partial_\mu \hX^I+B_\mu
\hX^I_+) (\partial_\nu \hX^J+B_\nu
\hX^J_+)\delta_{IJ}+
\nonumber \\
&+&(C_\mu^I-\partial_\mu
\hX_-^I)\delta_{IJ}
\partial_\nu \hX^J_+ \ , \quad I,J=1,\dots,8 \ . \nonumber \\
\end{eqnarray}
Then the action  up to quadratic approximation
can be written as
\begin{equation}
S=-\tau_{M2}\int d^3\xi
\sqrt{-\eta}-\tau_{M2}
\int d^3\xi \sqrt{-\det\eta}
\mL \ ,
\end{equation}
where the Lagrangian density takes the form
\begin{eqnarray}
\mL&=&\frac{1}{2} \eta^{\mu\nu}
(\partial_\mu \hX^I+B_\mu\hX^I_+)
\delta_{IJ} (\partial_\nu
\hX^J+B_\nu\hX^J_+)\delta_{IJ}
+\nonumber \\
&+&\frac{1}{2}
\eta^{\mu\nu}(C_\mu^I-\partial_\mu \hX^I_-)
\partial_\nu \hX^J_+\delta_{IJ}-
\nonumber \\
&-&\epsilon^{\mu\nu\rho}
(\pi\alpha' B_\mu \mF_{\nu\rho}+
m(\pi\alpha')^2V_\mu\partial_\nu V_\rho) \ .
\nonumber \\
\end{eqnarray}
that is again very close to the abelian
form of the action given in
\cite{Ezhuthachan:2008ch} and provides
further support of our construction.


\section{Conclusion}\label{fourth}
Let us summarize our results. We
studied EMP procedure for
Dirac-Born-Infeld action for D2-brane
and we found its limitation. Then we
suggested an alternative form of this
procedure that  is based on a
formulation of gauged M2-brane action.
This fact however implies that the
theory should be defined in background
with non-zero mass parameter $m$ and
this observation certainly deserves
better understanding and more detailed
 study. Further, it would be also
interesting to develop BRST Hamiltonian
treatment of the action (\ref{M2final})
and compare it with the similar
analysis that was given in
\cite{Benvenuti:2008bt}. Finally, it
will be extremely interesting to see
whether there exists an non-abelian
extension of the action
(\ref{M2final}). We hope to return to
these problems in future.
\\
\\
{\bf Acknowledgement}

This work
 was supported  by the Czech Ministry of
Education under Contract No. MSM
0021622409.

\begin{appendix}
\section{Direct Dimensional Reduction for
Massive M2-brane}\label{Appendix} In
this appendix we show that the gauged
M2-brane action upon direct dimensional
reduction in the direction $X$
associated to the gauged isometry
reduces to the action
(\ref{OrtinactDBIdr}). To begin with we
choose coordinates that are adapted to
the isometry so that
$\hk^M=\delta^{Mx}$ and we split eleven
coordinates $\hX^M$ into the ten
$10$-dimensional $X^m,m=0,\dots,9$ and
the extra scalar $\hX^{x}\equiv X$.
Using the relations between the
$11$-dimensional and $10$-dimensional
fields
\begin{eqnarray}\label{ansA}
\hg_{xx}=e^{\frac{4}{3}\Phi} \ , \quad
\hg_{mx}=e^{\frac{4}{3}\Phi} C^{(1)}_m
\ ,
\nonumber \\
\hg_{mn}=e^{-\frac{2}{3}\Phi}g_{mn}+
e^{\frac{4}{3}\Phi}C^{(1)}_m C^{(1)}_n
\ ,
\nonumber \\
\hC_{mnk}=C^{(3)}_{mnk} \ , \quad
\hC_{mnx}=B_{mn} \  \nonumber \\
\end{eqnarray}
 it is straightforward to see that
\begin{equation}\label{bAA}
\bA_{\mu\nu}=e^{-\frac{2}{3}\Phi}
\partial_\mu X^m\partial_\nu X^n g_{mn}+
e^{\frac{4}{3}\Phi}(\partial_\mu
X+B_\mu+C^{(1)}_\mu) (\partial_\nu
X+B_\nu +C^{(1)}_\nu) \ .
\end{equation}
Then if we insert (\ref{bAA}) together
with (\ref{ansA}) into the action
(\ref{M2gauged}) we easily obtain that
it  reduces to the  action
(\ref{OrtinactDBIdr}).
\end{appendix}

\end{document}